# Toward Quantum Enabled Solutions for Real-Time Currency Arbitrage in Financial Markets

### v1

@@ Abstract
Index Terms:
Currency Arbitrage, Quadratic Unconstrained Binary Optimization, Quantum Annealers, Adaptive Cost Encoding, Variational Quantum Algorithm.
@@ I Introduction
  • Quantum and classical solving: We solve this model using a range of solvers- quantum annealing, gate-based quantum hybrid method (ACE), and classical solvers (Gurobi and Tabu Search).
@@ III-A QUBO Formulation
$min\Bigg{(}\Bigg{.}\sum_{i,j\in V}\sum_{k=1}^{K}-log(w_{i,j})\cdot x_{i,k}x_{j,k+1}+\\ \lambda_{1}\sum_{k=1}^{K}(\sum_{i=1}^{N}x_{i,k}-1)^{2}+\lambda_{2}\sum_{i=1}^{N}\sum_{k=1,}^{K-1}\sum_{l=k+1}^{K}x_{i,k}x_{i,l}+\\ \lambda_{3}\sum_{i=1}^{N}\sum_{\begin{subarray}{c}j=1\\ (i,j)\notin E\end{subarray}}^{N}\sum_{k=1}^{K}x_{i,k}x_{j,k+1}+\lambda_{4}\sum_{i=1}^{N}(x_{i,1}-x_{i,K+1})^{2}\Bigg{.}\Bigg{)}$ (6)
@@ III-C Adaptive Cost Encoding
To tackle the challenge of solving large-scale QUBO problems on gate-based quantum computers, Adaptive Cost Encoding (ACE) algorithm leverages Variational Quantum Algorithms (VQAs) with a minimal encoding scheme. This approach exponentially reduces the qubit demands compared to traditional VQAs. For example, Variational Quantum Eigensolver (VQE) and Quantum Approximate Optimization Algorithm (QAOA), where required qubits are equal to number of binary variables in QUBO. In ACE to solve a QUBO problem with $n$ number of binary variables, the number of qubits reduce to $\left\lceil log_{2}(n)\right\rceil+1$. Here, $\left\lceil log_{2}(n)\right\rceil$ register qubits encode the state indices and an ancilla qubit represents the binary value corresponding to each state. The quantum circuit is prepared with a layer of Hadamard gates to create an equal superposition, subsequently a hardware-efficient ansatz is used to evolve the quantum state. Hardware-efficient ansatz includes sequence of parameterized single-qubit rotations and entangled gates. The quantum state has been provided by eq.(7).

### v2

@@ I Introduction
  • Quantum and classical solving: We solve this model using a range of solvers- quantum annealing, gate-based quantum hybrid method (ACE) bib7 , and classical solvers (Gurobi and Tabu Search).
@@ III-A QUBO Formulation
[…]
@@ III-C Adaptive Cost Encoding
To tackle the challenge of solving large-scale QUBO problems on gate-based quantum computers, Adaptive Cost Encoding (ACE) algorithm bib7 leverages Variational Quantum Algorithms (VQAs) with a minimal encoding scheme. This approach exponentially reduces the qubit demands compared to traditional VQAs. For example, Variational Quantum Eigensolver (VQE) and Quantum Approximate Optimization Algorithm (QAOA), where required qubits are equal to number of binary variables in QUBO. In ACE to solve a QUBO problem with $n$ number of binary variables, the number of qubits reduce to $\left\lceil log_{2}(n)\right\rceil+1$. Here, $\left\lceil log_{2}(n)\right\rceil$ register qubits encode the state indices and an ancilla qubit represents the binary value corresponding to each state. The quantum circuit is prepared with a layer of Hadamard gates to create an equal superposition, subsequently a hardware-efficient ansatz is used to evolve the quantum state. Hardware-efficient ansatz includes sequence of parameterized single-qubit rotations and entangled gates. The quantum state has been provided by eq.(7) bib7 .